# HiFo-Prompt: Prompting with Hindsight and Foresight for LLM-based Automatic Heuristic Design

### v1

@@ Abstract
LLM-based Automatic Heuristic Design (AHD) within Evolutionary Computation (EC) frameworks has shown promising results. However, its effectiveness is hindered by the use of static operators and the lack of knowledge accumulation mechanisms. We introduce HiFo-Prompt, a framework that guides LLMs with two synergistic prompting strategies: Foresight and Hindsight. Foresight-based prompts adaptively steer the search based on population dynamics, managing the exploration-exploitation trade-off. In addition, hindsight-based prompts mimic human expertise by distilling successful heuristics from past generations into fundamental, reusable design principles. This dual mechanism transforms transient discoveries into a persistent knowledge base, enabling the LLM to learn from its own experience. Empirical results demonstrate that HiFo-Prompt significantly outperforms state-of-the-art LLM-based AHD methods, generating higher-quality heuristics while achieving substantially faster convergence and superior query efficiency.
@@ 1 Introduction
Combinatorial Optimization (CO) is a significant branch of mathematical optimization (dreo2006metaheuristics, ), focusing on identifying solutions that maximize or minimize an objective function over a discrete feasible space.
Over the past decades, numerous efficient heuristic algorithms (BlumRoli2003, ; ColorniEtAl1996, ) have emerged in the field of optimization to solve these problems.
However, manual design of heuristic algorithms remains a task that heavily relies on domain expertise, requiring a deep understanding of problem structures and extensive experiential knowledge (CamachoVillalonEtAl2023, ).
In recent years, the rise of Large Language Models (LLMs) has brought new opportunities for the automated design of heuristic algorithms (VaswaniEtAl2017, ; WangChen2023, ; LiuYaoEtAl2024, ). A number of LLM-based Automated Heuristic Design (LLM-based AHD) methods have been proposed, many of which apply the Evolutionary Computation (EC) frameworks. Pioneering works such as FunSearch (romera2024, ) and EoH (liu_icml2024, ) have demonstrated the effectiveness of the evolutionary computation paradigm. FunSearch uses LLMs to evolve programs, searching for high-quality heuristic functions in a predefined heuristic program space. EoH introduces the co-evolution of thought and code, enabling LLMs to iterate between the thought process and code generation, thereby simulating the heuristic design process employed by human experts.
Following prior works, some LLM-based AHD approaches (vanSteinBack2025, ; ZhangLiuEtAl2024, ) focus on designing diverse guidance methods to enhance the search process. For example, ReEvo (ye2024, ) proposes a reflection-based framework to simulate language gradient guidance for directing the search process. HSEvo (Dat2025HSEvo, ) integrates diversity metrics with harmony search to balance population diversity and convergence better. However, these methods remain inadequate in terms of knowledge management wu2024 , as they do not incorporate a systematic mechanism for explicitly distilling and reusing algorithmic principles proven effective during evolution. Consequently, valuable insights generated by LLMs are often transient, being either discarded with the elimination of parent candidates or diluted over successive iterations.
Meanwhile, some LLM-based AHD approaches seek breakthroughs by refining the structure of the search process. MCTS-AHD (zheng2025, ) uses Monte Carlo Tree Search to manage heuristics, retaining temporarily underperforming but promising individuals to enable broader exploration. However, such methods typically embed regulation mechanisms passively within fixed algorithmic workflows, with knowledge utilization confined to local reasoning on individual population members. They lack an independent system capable of monitoring the global population’s state and proactively switching generation modes upon detecting issues.
In this paper, to address the limitations of the above approaches, we propose a novel framework named Hindsight-Foresight Prompt (HiFo-Prompt). At the individual level, we decouple thought from code and independently extract each individual’s thought, referred to as an insight. Then, these insights are added to an Insight Pool, a dynamic repository that accumulates and updates distilled knowledge during evolution. This process improves the framework’s Hindsight capability. At the population level, we introduce the Evolutionary Navigator, a control function that monitors the global state of the population after each iteration. Based on this state information, it adaptively determines the exploration–exploitation mode for the next iteration, strengthening the Foresight ability of the framework.
In summary, the contributions of our approach are as follows:
@@ 2 Related Work
2.1 LLM-based Automatic Heuristic Design
AHD (PotterCulleyDarlingtonChawdhry2003, ) has traditionally relied on Hyper-Heuristics (HH) for CO (drake2020, ; burke2009, ; DokerogluKucukyilmazTalbi2024, ). However, HH’s dependence on generating or selecting from handcrafted components (burke2009, ; burke2018, ) inherently constrains algorithmic novelty and adaptability (li2024, ). While the advent of LLMs opened new frontiers, initial work showed that naive prompting often fails to produce competitive heuristics for complex problems (yang2024, ; kristiadi2024, ; guo2024, ).
A more promising paradigm integrates LLM with EC, a synergy pioneered by Funsearch (romera2024, ). In this LLM with EC approach, the LLM acts as a sophisticated genetic operator (i.e., crossover and mutation). This method has fueled successful frameworks like EoH (liu_icml2024, ) and its successors, which have been rapidly extended to domains like code generation and heuristic design (ye2024, ; zheng2025, ; liu2023, ; zhang2025, ).
2.2 Knowledge Management in Generative Search
Leveraging historical information is critical for efficient search. In classical EC, Cultural Algorithms (maher2021comprehensive, ; reynolds2000cultural, ) employ a "Belief Space" to store and evolve collective knowledge, guiding population evolution in a structured manner (reynolds2000cultural, ). In the era of LLMs, this concept has been mirrored by "reflection" mechanisms, where models refine their outputs based on verbal feedback (shinn2023reflexion, ; bo2024reflectivemultiagent, ). However, this self-reflection often produces unstructured and volatile feedback that is difficult to reuse systematically across generations (zhang2024selfcontrast, ). Furthermore, it can be biased by the model’s pre-trained knowledge rather than objective empirical evidence from the ongoing search (zhang2024selfcontrast, ).
2.3 Adaptive Control in Evolutionary Computation
Dynamically adapting strategies based on the search state is a long-standing goal in EC. Traditional methods include fine-grained adjustments such as Adaptive Operator Selection (AOS) (fialho2010adaptive, ; Tian2023Deep, ) and parameter control (eiben2015, ; aleti2016systematic, ). Recently, researchers have begun to use LLMs’ reasoning capabilities for adaptive interventions, such as generating diverse individuals to escape local optima (yu2024deep, ; veerapen2012exploration, ). This marks a conceptual shift from reactive, low-level parameter tuning to proactive, high-level strategic reasoning, allowing the search to adapt based on a semantic understanding of the population’s state (Eiben1999, ; Karafotias2015, ; Papa2021Applications, ; deLacerda2021systematic, ). However, the efficacy of these interventions often relies on predefined triggers and static prompts, which may not be optimal across all stages of the evolutionary process (Zhang2012Survey, ; Kramer2010Evolutionary, ).
@@ 3 Methodology
In this section, we introduce our proposed HiFo-Prompt, a novel framework that equips the LLM-driven evolutionary process with mechanisms for learning and adaptation (shown in Figure 1). HiFo-Prompt integrates two synergistic components: a Foresight module for real-time adaptive control, which monitors evolutionary dynamics to steer the search strategy, and a Hindsight module for long-term knowledge accumulation, which manages a self-evolving repository of successful design principles that we term insights. By uniting foresight-driven strategy with hindsight-informed knowledge, our framework transforms the generative process into a robust, self-regulating system.
Figure 1: The framework of HiFo-Prompt. Left: Prompt Construction builds upon a foundational prompt, augmenting it with strategic Foresight and empirical Hindsight to form the final composite prompt. Right: Knowledge Evolution forms a learning loop where elite heuristics are distilled into new insights to continuously enrich the Hindsight knowledge base.
@@ 3.1 Guided Prompt Synthesis for AHD
Our HiFo-Prompt framework applies a Guided Prompt Synthesis mechanism that constructs each prompt as a context-aware composite instruction. This mechanism integrates three interlocking modules: Foundational prompt strategies, hint knowledge, and a foresight-driven strategy. The resulting composite prompt provides precise, multifaceted guidance to the LLM. A complete example of prompt generation for TSP with step-by-step construction is provided in Figure 2.
@@
Our framework’s generative foundation is a set of Foundational Prompt Strategies, which function as the LLM-equivalent of genetic operators. We first generate heuristics from scratch using the initial prompt strategy I1, then evolve them with five foundational prompts adapted from EoH liu_icml2024 . These prompts are organized into two primary strategies: 1) Reorganization Strategies, which include E1, synthesizing a new algorithm with a novel structure from multiple parents, and E2, abstracting shared core ideas to generate conceptually distinct variants; and 2) Mutation Strategies, which encompass M1, making structural modifications for functionally equivalent variants; M2, tuning critical parameters; and M3, simplifying components prone to overfitting. While this curated set provides the raw generative capability, its effectiveness depends on contextual guidance. This is the role of the Hindsight and Foresight modules, which inject insights and a design directive into the prompts to align each action with the search’s current needs.
@@
This module incorporates proven knowledge in the form of insights, which are abstract and generalizable design principles distilled from successful heuristics. These insights are managed in a dynamic Insight Pool, where each is assigned and continuously updated with a credibility score based on its empirical performance. Before generation, high-scoring insights are retrieved and embedded into the prompt. They serve as validated priors to steer the LLM toward promising designs. While effective for historical guidance, this module cannot address real-time evolutionary needs.
@@
The Foresight module implements real-time strategic control, orchestrated by its core component, the Evolutionary Navigator. The Navigator continuously monitors macroscopic evolutionary indicators, such as performance stagnation and population diversity, to assess the state of the search. Based on this analysis, it selects the governing evolutionary regime for the subsequent generation. This regime dictates the overall strategic focus by choosing one of three explicit modes (Crepinsek2013Exploration, ): 1) Explore, to foster novelty when diversity is low or progress has stalled; 2) Exploit, to refine high-performing solutions when progress is consistent; 3) Balance, to maintain a synergistic application of all operators. This high-level directive shapes the final prompt, ensuring that the generative process of the LLM aligns with the immediate strategic needs of the evolution.
@@ 3.2 Hindsight: Mechanisms of the Self-Evolving Insight Pool
While foundational prompts provide the generative actions, the Hindsight module steers evolution with empirically validated knowledge managed within the Insight Pool. Analogous to the Belief Space in Cultural Algorithms (coello2010efficient, ), our work introduces a novel instantiation for LLM-based AHD, distinguished by two core features: LLM-driven insight extraction and a utility-based credit assignment tailored for prompt-based generation.
To mitigate the cold-start problem, we initialize the Insight Pool with seed insights from established heuristic design literature. The core objective of the framework then becomes the autonomous discovery and refinement of new, problem-specific insights. This is achieved through a continuous lifecycle that systematically transforms transient evolutionary successes into explicit, reusable knowledge assets, governed by three integrated phases: insight extraction, utility-driven application, and adaptive pruning.
@@
The lifecycle begins by expanding the knowledge base. At the end of each generation, we prompt an LLM to distill generalizable design principles from the elite individuals of the population (see Appendix G.1). To preserve informational diversity, these candidate insights are admitted only if their Jaccard similarity to all existing pool members falls below a novelty threshold, thus preventing redundancy. Admitted insights become active candidates for guiding future generations.
@@
To guide heuristic generation, we employ a utility-based retrieval mechanism. Instead of random sampling, it selects the top-$s$ insights with the highest adaptive utility score $U(k_{i},t)$. Subsequently, we use a credit assignment (Whitacre2006Credit, ) to update these utility scores, rewarding insights that lead to high-performing heuristics. The utility function is formulated to balance exploitation and exploration:
$U(k_{i},t)=\underbrace{E_{i}(t)}_{\text{Effectiveness}}-\underbrace{w_{u}\log(N_{i}(t)+1)}_{\text{Usage Penalty}}+\underbrace{B_{r}(t,\tau_{i})}_{\text{Recency Bonus}}$ (1)
where $E_{i}(t)$ is the learned effectiveness, the penalty term (weighted by $w_{u}$) discourages overuse, and the recency bonus $B_{r}$ promotes strategic coherence by rewarding insights used in recent successful generations.
The effectiveness score $E_{i}(t)$ is updated via credit assignment. This process first converts the raw fitness $g(h_{\text{new}})$ of an offspring into a normalized, problem-agnostic score $\tilde{\rho}$, scaling its performance relative to the current population’s best and worst solutions:
$\tilde{\rho}=\frac{g(h_{\text{worst}})-g(h_{\text{new}})}{g(h_{\text{worst}})-g(h_{\text{best}})}$ (2)
We posit that an offspring’s evolutionary contribution is non-linear. Therefore, we map $\tilde{\rho}$ to a final credit signal, $g_{\text{eff}}$, using a tiered, piecewise function. This deliberate design creates distinct reward regimes for qualitatively different outcomes:
$g_{\text{eff}}=\begin{cases}0.8+0.2\cdot\tilde{\rho}&\text{if }g(h_{\text{new}})\leq g(h_{\text{best}})\\ 0.2+0.6\cdot\tilde{\rho}&\text{if }g(h_{\text{best}})<g(h_{\text{new}})\leq g(h_{\text{avg}})\\ -0.3+0.5\cdot\tilde{\rho}&\text{if }g(h_{\text{new}})>g(h_{\text{avg}})\end{cases}$ (3)
This structure provides strong positive signals for paradigm shifts, moderate rewards for incremental progress, and penalties for below-average performance. Finally, for each insight $j$ in the contributing set $K_{c}$, its effectiveness score is updated through an Exponential Moving Average (EMA):
@@
where the learning rate $\alpha\in[0,1]$ is set to a small value to smooth the stochastic credit signals from individual evaluations. This allows an insight’s robust, long-term utility to emerge statistically, preventing its score from being skewed by outlier performances.
@@
To maintain quality within its finite capacity $C_{\text{pool}}$, the pool employs an adaptive pruning mechanism triggered when $|K_{\text{pool}}|>C_{\text{pool}}$. This process removes the insight with the minimum eviction score, $\mathcal{S}_{\text{evict}}$, which balances an insight’s proven performance against the risk of its obsolescence:
@@
where $E_{i}(t)$ is the current effectiveness of the insight, $t_{\text{last\_used}}(k_{i})$ is the generation of its last use, and the time decay rate $R_{\text{decay}}$ targets not only low-effectiveness insights, but also those that have become inactive. The decay rate is set conservatively to prevent the premature removal of valuable but temporarily dormant knowledge. Furthermore, a grace period grants new insights with usage counts below a threshold $T_{\text{usage}}$, eviction immunity. This ensures novel principles are given a fair opportunity to demonstrate their utility before facing pruning.

### v2

@@ Abstract
This paper investigates the application of Large Language Models (LLMs) in Automated Heuristic Design (AHD), where their integration into evolutionary frameworks reveals a significant gap in global control and long-term learning. We propose the Hindsight-Foresight Prompt (HiFo-Prompt), a novel framework for LLM-based AHD designed to overcome these limitations. This is achieved through two synergistic strategies: Foresight and Hindsight. Foresight acts as a high-level meta-controller, monitoring population dynamics(e.g., stagnation and diversity collapse) to switch the global search strategy between exploration and exploitation explicitly. Hindsight builds a persistent knowledge base by distilling successful design principles from past generations, making this knowledge reusable. This dual mechanism ensures that the LLM is not just a passive operator but an active reasoner, guided by a global plan (Foresight) while continuously improving from its cumulative experience (Hindsight). Empirical results demonstrate that HiFo-Prompt significantly outperforms a comprehensive suite of state-of-the-art AHD methods, discovering higher-quality heuristics with substantially improved convergence speed and query efficiency. Our code is available at https://github.com/Challenger-XJTU/HiFo-Prompt.
@@ 1 Introduction
Combinatorial Optimization (CO) problems, which involve finding an optimal solution from a discrete set of possibilities, are ubiquitous in science and engineering. Because of their NP-hardness, designing effective heuristics for these problems is a complex task, traditionally based on extensive human experience and intuition (Camacho-Villalón et al., 2023).
The advent of Large Language Models (LLMs) has catalyzed a paradigm shift toward Automated Heuristic Design (AHD) (Wang and Chen, 2023; Liu et al., 2024c). A particularly potent approach marries LLMs with Evolutionary Computation (EC), casting the LLM as a high-level semantic mutation operator. Foundational works such as FunSearch (Romera-Paredes et al., 2024) and EoH (Liu et al., 2024b) established the viability of this LLM+EC paradigm, demonstrating its capacity to discover novel and effective heuristics.
However, as the field progresses, two fundamental challenges have emerged in AHD: the inability to steer the heuristic generation process based on population dynamics and the failure to distill and manage the core design principles of high-performance heuristics to guide the subsequent heuristic generation process.
First, many approaches lack a mechanism for global adaptive guidance. They often rely on local or reactive signals; for instance, ReEvo (Ye et al., 2024) performs reflection on a single candidate, while methods such as MCTS-AHD (Zheng et al., 2025) passively embed the exploration-exploitation trade-off within their search structure. This localized control does not respond to the macroscopical dynamics of the population and cannot proactively intervene when the search encounters systemic issues such as premature convergence or a decline in diversity. A more aggressive strategy involves in-weight adaptation (e.g., EvoTune (Šurina et al., 2025), CALM (Huang et al., 2025)), which uses numerical gradients to fine-tune the LLM. Although powerful, this approach incurs high computational costs from repeated fine-tuning and reduces the LLM to an opaque policy network. Consequently, learning is implicitly encoded in model weights, preventing the extraction of explicit design principles and obscuring the model’s symbolic reasoning process.
Second, existing frameworks suffer from poor knowledge persistence, a phenomenon we term knowledge decay. Successful design strategies often remain entangled within specific code implementations; when parent candidates are discarded, the underlying logic is lost. Recent advances, such as evolving the optimizer in MoH (Shi et al., 2025) or automating problem reduction in RedAHD (Thach et al., 2025), operate in orthogonal dimensions. They do not explicitly decouple algorithmic knowledge from executable forms. Consequently, the system fails to achieve cumulative learning, perpetually rediscovering similar concepts instead of building on proven algorithmic strategies.
To overcome these fundamental limitations, we propose HiFo-Prompt (Hindsight-Foresight Prompt), a framework that establishes a hierarchical control architecture for LLM-based AHD. HiFo-Prompt elevates the LLM from a mere code generator to a symbolic meta-optimizer by endowing it with two synergistic capabilities: First, the Foresight module addresses the control problem by serving as a meta-controller(Evolutionary Navigator) that observes population dynamics. Upon detecting states like performance plateaus, it explicitly modulates the generative process by switching evolutionary regimes via transparent verbal gradients injected at the prompt level, serving as a symbolic alternative to opaque and expensive numerical gradients. Second, the Hindsight module addresses knowledge decay by implementing the Insight Pool, an evolving repository of abstracted design principles. It distills core algorithmic patterns from specific code implementations, transforming them into reusable knowledge. This mechanism allows the system to build on validated design principles, effectively seeding subsequent generations with proven algorithmic strategies. In summary, the contributions of our approach are as follows:
@@ 2 Related Work
LLM-driven Automatic Heuristic Design
The integration of Large Language Models (LLMs) into Evolutionary Computation (EC) is a vibrant new direction for Automated Heuristic Design (AHD) (Liu et al., 2024a; Chauhan et al., 2025). Pioneered by works like FunSearch (Romera-Paredes et al., 2024) and EoH (Liu et al., 2024b), this paradigm leverages the LLM as a powerful semantic operator to generate heuristics as code. Recent efforts to advance this paradigm can be categorized along several axes. Some works focus on refining search control through sophisticated prompt engineering and guidance mechanisms (Ye et al., 2024; Dat et al., 2025), or by redesigning the population structure itself (Zheng et al., 2025). A distinct, more model-centric approach directly adapts the LLM’s parameters via reinforcement learning-based fine-tuning (Šurina et al., 2025; Huang et al., 2025). At the highest level of abstraction, research has also explored evolving other core components of the optimization process, such as the optimizer (Shi et al., 2025) or the problem representation (Thach et al., 2025).
Knowledge Management in Generative Search
Harnessing historical information is a cornerstone of efficient search. In classical EC, methods like Cultural Algorithms (Maheri et al., 2021) formalize this via a structured Belief Space that stores and evolves collective knowledge. Contemporary LLM-based approaches often rely on in-context reflection mechanisms (Shinn et al., 2023; Bo et al., 2024), where the model self-critiques failures to inform its next attempt. However, this knowledge is typically transient, unstructured, and instance-specific. Consequently, these methods lack a mechanism for accumulating and generalizing insights over time, preventing the formation of a persistent, structured knowledge base analogous to those in classical EC.
Adaptive Control in Evolutionary Computation
Dynamically adapting search strategies is a long-standing goal in EC. Historically, this has been addressed through low-level, reactive mechanisms like Adaptive Operator Selection (AOS) (Fialho, 2010; Tian et al., 2023) and parameter control (Eiben and Smith, 2015; Aleti and Moser, 2016). These methods rely on numerical credit assignment, effectively voting for strategies that recently performed well, but they lack a semantic understanding of the search dynamics (e.g., identifying population stagnation). The reasoning capabilities of LLMs offer a paradigm shift towards higher-level, proactive control (Eiben et al., 1999; Papa, 2021). Instead of merely adjusting parameters, LLMs can interpret population-level statistics to suggest symbolic actions, such as increase mutation rate to escape a local optimum. This marks a transition from fine-grained numerical tuning to semantic-based strategic adjustment.
@@ 3 Methodology
Figure 1: The framework of HiFo-Prompt, which comprises two core processes. (1) Prompt Construction: A foundational prompt is dynamically augmented with design directives from the Foresight and insights from the Hindsight to form the final composite prompt. (2) Knowledge Evolution: A self-evolving loop is established where elite heuristics from the evolutionary process are distilled into new insights, continuously enriching the Hindsight module’s knowledge base.
In this section, we introduce our proposed HiFo-Prompt, a novel framework that equips the evolutionary process driven by LLM with mechanisms for learning and adaptation (shown in Figure 1). HiFo-Prompt integrates two synergistic components: a Foresight module for real-time adaptive control, which monitors evolutionary dynamics to steer the search strategy, and a Hindsight module for long-term knowledge accumulation, which manages a self-evolving repository of successful design principles that we term insights. By uniting foresight-driven strategy with hindsight-informed knowledge, our framework transforms the generative process into a robust, self-regulating system.
@@ 3.1 Guided Prompt Synthesis for AHD
Our HiFo-Prompt framework applies a Guided Prompt Synthesis mechanism that constructs each prompt as a context-aware composite instruction. This mechanism integrates three interlocking modules: Foundational prompt strategies, Hindsight, and Foresight. The resulting composite prompt provides precise, multifaceted guidance to the LLM. A complete example of prompt generation for TSP with step-by-step construction is provided in Figure 2, and the specific templates for all operators (e.g., I1, E1, M1) are detailed in Appendix E.
@@
Our framework’s generative foundation is a set of Foundational Prompt Strategies, which function as the LLM-equivalent of genetic operators. We first generate heuristics from scratch using the initial prompt strategy I1, then evolve them with five foundational prompts adapted from EoH  (Liu et al., 2024b). These prompts are organized into two primary strategies: 1) Reorganization Strategies, which include E1, synthesizing a new algorithm with a novel structure from multiple parents, and E2, abstracting shared core ideas to generate conceptually distinct variants; and 2) Mutation Strategies, which encompass M1, making structural modifications for functionally equivalent variants; M2, tuning critical parameters; and M3, simplifying components prone to overfitting. While this curated set provides the raw generative capability, its effectiveness depends on contextual guidance. This is the role of the Hindsight and Foresight modules, which inject insights and a design directive into the prompts to align each action with the search’s current needs.
@@
This module incorporates valid heuristic experience in the form of insights, which are abstract and generalizable design principles distilled from successful heuristics. These insights are managed in a dynamic Insight Pool, where each is assigned and continuously updated with a credibility score based on its empirical performance. Before generation, high-scoring insights are retrieved and embedded into the prompt. They serve as validated priors to steer the LLM toward promising designs. While effective for historical guidance, this module cannot address real-time evolutionary needs.
@@
The Foresight module implements real-time evolutionary state control, orchestrated by its core component, the Evolutionary Navigator. The Navigator continuously monitors macroscopic evolutionary indicators, such as performance stagnation and population diversity, to assess the state of the search. Based on this analysis, it selects the governing evolutionary regime for the subsequent generation. This regime dictates the evolutionary search direction by choosing one of three explicit modes (Črepinšek et al., 2013): 1) Explore, to foster novelty when diversity is low or progress has stalled; 2) Exploit, to refine high-performing solutions when progress is consistent; 3) Balance, to maintain a synergistic application of all operators. This directive guides the prompt, ensuring the LLM adapts to the current search phase.
@@ 3.2 Hindsight: Mechanisms of the Self-Evolving Insight Pool
Initially, HiFo-Prompt initializes the Insight Pool with seed insights(See Section F.3) to bootstrap the generative process. Serving as an initial scaffold, these seeds allow the framework to immediately shift focus to autonomously evolving novel heuristics. This is achieved through a continuous lifecycle that systematically transforms transient evolutionary successes into reusable knowledge assets, governed by three integrated phases: insight extraction, utility-driven application, and adaptive pruning.
Initially, HiFo-Prompt initializes the Insight Pool with seed insights(See Section F.3) to bootstrap the generative process. Serving as an initial scaffold, these seeds allow the framework to immediately shift focus to autonomously evolving novel, problem-specific heuristics. This is achieved through a continuous lifecycle that systematically transforms transient evolutionary successes into explicit, reusable knowledge assets, governed by three integrated phases: insight extraction, utility-driven application, and adaptive pruning.
@@
The lifecycle begins by expanding the knowledge base. At the end of each generation, we prompt an LLM to distill generalizable design principles (insights) from the elite individuals of the population (See Section F.1). To preserve informational diversity, a candidate insight $k_{\text{new}}$ is admitted to the insight pool $K_{\text{pool}}$ only if its Jaccard similarity to all existing pool members falls below a novelty threshold $\theta_{\text{novelty}}$. For this comparison, each insight’s text is preprocessed by converting it to lowercase and tokenizing it based on whitespace. The insights themselves become active candidates for the guidance of future generations.
@@
To guide heuristic generation, we employ a utility-based retrieval mechanism. For each new generation attempt, the mechanism selects the top-$s$ insights with the highest adaptive utility score $U(k_{i},t)$. This contributing set of insights, denoted as $K_{c}$, is then injected into the LLM’s prompt. Following the evaluation of the offspring generated, we use credit assignment (Whitacre et al., 2006) to update the utility scores of all the insights in $K_{c}$. The utility function is formulated to balance exploitation and exploration:
$U(k_{i},t)=\underbrace{E_{i}(t)}_{\text{Effectiveness}}-\underbrace{w_{u}\log(N_{i}(t)+1)}_{\text{Usage Penalty}}+\underbrace{B_{r}(t,t_{i}^{\text{last}})}_{\text{Recency Bonus}}$ (1)
where $E_{i}(t)$ is the learned effectiveness of insight $i$. The penalty term, weighted by $w_{u}$, discourages overuse by penalizing an insight based on its total retrieval count $N_{i}(t)$, thus promoting exploration of less-used ideas. The recency bonus $B_{r}$ offers a temporary reward for insights used recently; specifically, it grants a fixed bonus if the insight was used within a small generation window $T_{w}$ (that is, if $t-t_{i}^{\text{last}}\leq T_{w}$), promoting strategic coherence.
The effectiveness score $E_{i}(t)$ is updated via credit assignment. This process first converts the raw fitness $g(h_{\text{new}})$ of an offspring into a normalized, problem-agnostic score $\tilde{\rho}$, scaling its performance relative to the best ($g(h_{\text{best}})$) and worst ($g(h_{\text{worst}})$) solutions of the current population:
$\tilde{\rho}=\frac{g(h_{\text{worst}})-g(h_{\text{new}})}{g(h_{\text{worst}})-g(h_{\text{best}})+\epsilon}$ (2)
where $\epsilon$ is a small constant to prevent division by zero. Drawing upon techniques for handling sparse rewards in Hierarchical Reinforcement Learning, we posit that an offspring’s evolutionary contribution is non-linear. Therefore, we map $\tilde{\rho}$ to a final credit signal, $g_{\text{eff}}$, using a tiered, piecewise function. This design creates distinct reward regimes for qualitatively different outcomes:
$g_{\text{eff}}^{\text{raw}}=\begin{cases}0.8+0.2\cdot\tilde{\rho}&\text{if }g(h_{\text{new}})\geq g(h_{\text{best}})\\ 0.2+0.6\cdot\tilde{\rho}&\text{if }g(h_{\text{best}})>g(h_{\text{new}})\geq g(h_{\text{avg}})\\ -0.3+0.5\cdot\tilde{\rho}&\text{if }g(h_{\text{new}})<g(h_{\text{avg}})\end{cases}$ (3)
This structure provides strong positive signals for paradigm-changing improvements ($\tilde{\rho}\geq 1$), moderate rewards for incremental progress ($0\leq\tilde{\rho}<1$), and penalties for below-average performance. The allowance of negative credit is a deliberate design choice to accelerate the pruning of detrimental ideas.
This piecewise reward structure incorporates techniques for handling sparse rewards in Hierarchical Reinforcement Learning. It provides discrete and explicit signals for distinct performance levels to enhance interpretability. The coefficients follow a specific asymmetric philosophy where the reward for elite solutions is significantly larger than the penalty magnitude, which is in turn larger than the reward for incremental improvements. This prioritization provides a strong signal for paradigm-shifting breakthroughs while effectively pruning unproductive directions. Furthermore, the parameter sensitivity analysis detailed in Appendix C.7 demonstrates that the framework maintains robust performance across a range of coefficient settings. This confirms that the structural logic of the reward mechanism is more critical than precise tuning.
To ensure stable updates, the raw credit is then clipped to a final value $g_{\text{eff}}=\max(-1.0,\min(1.0,g_{\text{eff}}^{\text{raw}}))$. Finally, for each insight $j$ in the contributing set $K_{c}$, its effectiveness score is updated through an Exponential Moving Average (EMA):
@@
where the learning rate $\alpha\in[0,1]$ smooths the stochastic credit signals from individual evaluations. This allows an insight’s long-term utility to emerge statistically, preventing its score from being skewed by outlier performances.
@@
To maintain quality within its finite capacity $C_{\text{pool}}$, the pool employs an adaptive pruning mechanism triggered when $|K_{\text{pool}}|>C_{\text{pool}}$. This process removes the insight with the minimum eviction score, $\mathcal{S}_{\text{evict}}$, which balances the proven performance of an insight against the risk of its obsolescence:
@@
where $E_{i}(t)$ is the current effectiveness of the insight, $t_{\text{last\_used}}(k_{i})$ is the generation of its last use, and the time decay rate $R_{\text{decay}}$ targets not only low-effectiveness insights, but also those that have become inactive. The decay rate is set conservatively to prevent the premature removal of valuable but temporarily dormant knowledge. Additionally, to prevent the premature elimination of underexplored principles, insights with usage counts below a threshold $T_{\text{usage}}$ are exempt from pruning, allowing sufficient iterations to establish their true utility.